Estimate each of the following using general rule: 
 12904 + 2888 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12904 + 2888
Round off to thousands
12904 is rounds off to 13000
2888 is rounds off to 3000
Hence, 12904 + 2888 = 13000 + 3000 = 16000
​​
Thanks